OpenStudy (anonymous):

Express 5^3 = x as a logarithmic equation.

OpenStudy (anonymous):

5^3=x

OpenStudy (mathmale):

Here you have an expression in the form x= (base 5)^y. In other words, 5 is the base of your exponential function here. Take the log to the base 5 of both sides of the given expression to convert it into a logarithmic equation:\[\log_{5} (5^3)=\log_{5} x\]

OpenStudy (mathmale):

Note that the exponential and logarithmic functions are inverses of one another. Use this fact to simplify the left side of the above equation. Where have you seen expressions of the form\[\log_{a} a^n~before?\]

OpenStudy (mathmale):

How have you simplified them?

OpenStudy (anonymous):

log_5 125=x

OpenStudy (anonymous):

\[\log_{5} 125=x\]

OpenStudy (mathmale):

Actually, \[\log_{a} a^n= n\] In words, the log operator cancels out the work of the exponential function, leaving us with simply n (the exponent). \[\log_{5} 5^4=?\]

OpenStudy (anonymous):

i'm confused

OpenStudy (mathmale):

What info do you need? What would resolve your confusion? If you're working with exponential and logarithmic functions, then you have seen the topic "inverse functions" before.

OpenStudy (anonymous):

not really i'm trying to learn it now